A Scientist has one solution that is 10% iodine and another that is 40% iodine. How much of each should the scientist use to produce 100mL of a mixture that is 25% iodine.
Answer by checkley77(12844)

.10x*.40(100-x)=.25*100
.10x+40-.40x=25
-.30x=25-40
-.30x=-15
x=-15/-.30
x=50 ml. of 10% iodine is used.
100-50=50 ml. of 40% iodine is used.
Proof:
.10*50+.40*50=.25*100
5+20=25
25=25